# S2Edit: Text-Guided Image Editing with Precise Semantic and Spatial Control

### v1

@@ 1 Introduction
Recent years have witnessed remarkable progress in the field of generative models, with techniques such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Karras et al., 2019, 2020, 2021) and Diffusion Models (Sohl-Dickstein et al., 2015; Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022) capable of synthesizing realistic images. These advancements have fueled significant interest in the realm of image editing, a highly practical and versatile domain. Among its various settings, text-guided editing stands out for its user-friendly interface, as it accepts natural language as the input.
@@
Early methods in this field often train GANs from scratch on paired image-text data (Dong et al., 2017; Li et al., 2020; Nam et al., 2018), requiring expensive annotations. Later approaches explore the latent space of pre-trained models like CLIP (Radford et al., 2021), which align text and image features. However, GANs are still trained on relatively small datasets, resulting in a latent space with limited capacity. They sometimes struggle to represent real-world objects at the tails of distributions faithfully, and fail to generalize to unseen text prompts. Recent advances in large-scale pre-trained text-to-image diffusion models (Saharia et al., 2022; Rombach et al., 2022) offer alternative solutions to this task. Thanks to the large scale pre-trained data, these methods exhibit a significantly stronger capacity in representing diverse objects from different domains, and are generally more robust to text prompt variations. Nevertheless, demonstrations of these editing methods’ performance typically involve marked changes in art styles, object shapes, motions, and precise editing of real-world images such as faces remains under-explored. When guided by detailed prompts, these methods frequently introduce severe artifacts or unrealistic styles. Moreover, they sometimes fail to edit the image utterly, with object identity information changed or the target attribute indicated in the prompt missing in the output. Essentially, current methods struggle to balance between identity information preservation and target prompt alignment (Zhang et al., 2023; Kawar et al., 2023; Brooks et al., 2023), and significant challenges still remain in editing real-world images both accurately and faithfully.
@@
In response to these challenges, we propose S2Edit, a two-stage text-guided image editing method. The first stage involves fine-tuning a pre-trained diffusion model to reconstruct the input image conditioned on a text prompt, which helps capture the object identity. To apply target edits while keeping the original identity information intact, we insert a learnable identity token into the prompt to capture nuanced identity features. However, plain fine-tuning lacks the control of the semantic meaning or spatial emphasis of the identity token, which is critical for precise image editing. We observe that the learned identity token often contains information on attributes targeted for modification, which impedes the later editing process. To disentangle the identity token from target attributes, we apply semantic control by enforcing an orthogonality constraint between the identity token and the prompt in the textual feature space during fine-tuning. Moreover, the identity token may learn to represent identity-irrelevant image regions, undermining the preservation of identity information. Therefore, we apply spatial control during fine-tuning by manipulating the cross-attention maps of the identity token with an object mask. In the second stage of S2Edit, we augment the cross-attention injection strategy from Prompt-to-Prompt (Hertz et al., 2022) with spatial control, and employ it for editing with our identity token. With its high-quality editing results, S2Edit outperforms state-of-the-art methods both quantitatively and qualitatively in evaluations, and garners at least 45% more preference from the participants in a user study.
@@ 2 Related Work
Diffusion-based Image Editing. First introduced by Sohl-Dickstein et al. (Sohl-Dickstein et al., 2015), diffusion models have recently achieved state-of-the-art image synthesis quality (Ho et al., 2020; Dhariwal & Nichol, 2021; Nichol & Dhariwal, 2021; Saharia et al., 2022; Rombach et al., 2022). Based on these models, image editing works perform tasks like inpainting (Song et al., 2020; Yang et al., 2023), stroke-based editing (Meng et al., 2021; Avrahami et al., 2022, 2023), and text-guided editing (Hertz et al., 2022; Tumanyan et al., 2023; Parmar et al., 2023). In particular, text-guided editing garners considerable interest due to its friendly interface and substantial control capabilities. Studies in this domain enable rich text guidance by manipulating attention maps (Hertz et al., 2022; Tumanyan et al., 2023; Parmar et al., 2023), text prompts (Nichol et al., 2022; Mokady et al., 2023; Brooks et al., 2023; Kawar et al., 2023; Pan et al., 2023; Dong et al., 2023), and feature space (Kim et al., 2022; Bansal et al., 2023). Prompt-to-Prompt (Hertz et al., 2022) investigates the relation between prompt tokens and spatial layout of images, and edits images by manipulating the cross-attention layers. InstructPix2Pix (Brooks et al., 2023) fine-tunes a pre-trained model on synthesized paired images and editing instructions. Despite impressive results, Prompt-to-Prompt sometimes fails to balance identity preservation and desired editing effects, and its performance often depends on the quality of the given prompt. InstructPix2Pix partially addresses the second issue with more straightforward instruction guidance, yet the first issue remains as it does not provide identity-related control.
@@
Personalized Image Synthesis and Manipulation. To preserve the fidelity of the original image, recent works in the image diffusion domain propose to insert learnable tokens and fine-tune pre-trained text-to-image models such that identity information can be learned and re-combined with new contexts (Gal et al., 2022; Kumari et al., 2023; Ruiz et al., 2023; Shi et al., 2023). DreamBooth (Ruiz et al., 2023) learns to encode object identity in a unique token by fine-tuning diffusion models on several images containing the same object. SINE (Zhang et al., 2023) investigates the overfitting problem when fine-tuning involves only one image, and proposes to distill the knowledge of the fine-tuned model into the original model. Instead of fine-tuning the entire model, Custom Diffusion (Kumari et al., 2023) only optimizes the Key and Value projection layers in cross-attention, and is able to compose multiple concepts via joint training on images with various objects. Although personalized fine-tuning helps preserve object identity, we observe a lack of semantic and spatial control over the learned identity token in existing methods, leading to failure modes such as entangled concepts when applied to detailed editing tasks.
@@
GAN-based Image Editing. One line of work on GAN attempts to discover meaningful feature space directions (Härkönen et al., 2020; Shen et al., 2020; Chen et al., 2022; Wu et al., 2021) that correspond to different semantic changes of images. Some other methods invert images to the latent space of pre-trained GANs (Richardson et al., 2021; Tov et al., 2021; Wang et al., 2022; Lyu et al., 2023; Wang et al., 2023) and leverage the aforementioned semantic directions for editing. This direction is further spurred by the development of large vision-language models such as CLIP (Radford et al., 2021). DeltaEdit (Lyu et al., 2023) proposes a CLIP delta space where visual and textual features from CLIP are better aligned, and trains a mapping network that connects the delta space to StyleGAN’s (Karras et al., 2020) latent space to enable text-guided editing. However, compared to state-of-the-art diffusion models, GANs, which are trained on smaller datasets, have a more limited expressibility in their latent space.
@@ 3 Method
Figure 2: S2Edit overview. Left: Given a source image and a text prompt, we insert a learnable token [I] into the text prompt and fine-tune a pre-trained text-to-image diffusion model to learn the identity information. To obtain a disentangled identity token, we apply an orthogonality constraint in the text embedding space via a semantic loss $L_{semantic}$ and force [I] to only represent the object of interest with masked cross-attention. Right: With [I] learned, we freeze the fine-tuned model and perform Null-text Inversion (Mokady et al., 2023) to get an initial noise map, then denoise it conditioned on the target prompt to generate the editing result.
@@ 3.1 Identity Fine-tuning
Instead of optimizing the prompt in the text space, we adapt the special token introduced in DreamBooth (Ruiz et al., 2023) to our method by constructing enhanced prompts $\widetilde{\mathcal{P}}$ and $\widetilde{\mathcal{P}^{\ast}}$ with a unique identity token [I] (e.g. “a [I] man with beard”), which encodes rich identity information and provides clear guidance to the generation process. Different from DreamBooth, which aims to extract appearance of a subject from multiple images for recontextualization tasks, we learn the identity information from one image to ensure that subtle details can be flawlessly preserved. We fine-tune the diffusion model and its text encoder with $\mathcal{I}$ and the enhanced prompt $\widetilde{\mathcal{P}}$ to learn the identity information, which is encoded into both the embedding of [I] and the model weights.
@@ 3.2 Precise Semantic and Spatial Control
where $\mathcal{A}_{\text{[I]}}$ is the original query-key cross-attention map of [I], $\mathcal{A}^{\ast}_{\text{[I]}}$ is its masked version, $M_{obj}$ is the binary mask of the object of interest, and $\odot$ is element-wise product. To obtain $M_{obj}$, we take the cross-attention map of the word that describes the object of interest (e.g., “lady”, “cat”) and binarize it, thus requiring no mask annotation. Compared to other text-guided image editing techniques, such as DeltaEdit, Imagic, SINE, and Prompt Tuning Inversion (Lyu et al., 2023; Kawar et al., 2023; Zhang et al., 2023; Dong et al., 2023), our approach imposes no additional demands on users, apart from specifying the object within the source prompt.
@@ 4.1 Experimental Setup
Implementation Details. While S2Edit can be applied to any text-to-image diffusion model, we implement it based on Stable Diffusion v1-4 (Rombach et al., 2022) due to its public availability and importance in the literature. During identity fine-tuning, we use the AdamW (Loshchilov & Hutter, 2017) optimizer with a base learning rate of $2\times 10^{-6}$. The semantic loss weight $\lambda$ is set to $0.1$. We fine-tune the model for 200 steps to avoid overfitting. For the DDIM scheduler, the number of diffusion steps $T$ is set to 50, and the other hyperparameters are the same as Stable Diffusion (Rombach et al., 2022). All experiments are conducted on one NVIDIA A100 GPU, where fine-tuning takes around 95 seconds, Null-text inversion process takes 113 seconds, and the inference time is 9 seconds. For the same source prompt and input image, our method requires fine-tuning and inversion to be run once, then only needs 9 seconds to generate edited results with different target prompts.
Datasets. To verify the editing performance and generalization ability of our proposed method, we conduct extensive experiments on images of diverse objects. Given the higher sensitivity of humans to detailed alterations and unnatural features in human faces, face editing serves as an ideal experiment for editing performance evaluation. Hence, we categorize source images into face and non-face domains, showcasing more results in face editing following (Lyu et al., 2023). For the human face domain, we use images from FFHQ (Karras et al., 2019) and CelebA (Liu et al., 2015). For non-face domains, we provide results on AFHQ (Choi et al., 2020), LSUN (Yu et al., 2015) cat and church images.
Comparisons. We compare S2Edit with several state-of-the-art diffusion-based text-guided image editing methods: Null-text Inversion (Mokady et al., 2023) combined with Prompt-to-Prompt (Hertz et al., 2022), InstructPix2Pix (Brooks et al., 2023), and SINE (Zhang et al., 2023). InstructPix2Pix accepts editing instructions as its prompt inputs, such as “Make her smile”. In addition, we compare with a state-of-the-art GAN-based method DeltaEdit (Lyu et al., 2023), which takes prompts inputs containing the word “face”, e.g., “Face with smile”.
@@ 4.2 Qualitative Evaluation
Fig. 3 shows the qualitative comparison between our method and four state-of-the-art methods on face editing. Our method faithfully preserves the original identities while accurately performing the desired edits. DeltaEdit (Lyu et al., 2023) follows the editing prompts well, but fails to preserve the identity of source images due to the limited capacity of GANs’ latent space. For example, the hairstyle is changed in the top row, and the expression drifts in the fourth row. InsturctPix2Pix (Brooks et al., 2023) makes the edits but fails to locate the desired editing area, e.g., removing only part of the glasses in the third row, and colorizing the glasses which are undesired in the last two rows. Except for the eyeglass removal failure, Null-text Inversion (Mokady et al., 2023) struggles to balance identity preservation and prompt alignment. For instance, the glasses are not removed in the third row, and identity drifts in the bottom row. SINE (Zhang et al., 2023) mostly produces unrealistic results with severe artifacts. In Fig. 4, we experiment with different edits for the same human, showing versatile editing capability of S2Edit. Only the specified attribute of interest is manipulated, while irrelevant features are unaltered, which proves that S2Edit learns disentangled identity tokens. Overall, S2Edit achieves the most accurate editing results on human faces which requires precise control of localized details.
@@ 4.3 Quantitative Evaluation
We report common metrics of all methods including FID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), and PSNR in Tab. 1. LPIPS and PSNR measure identity preservation, and FID reflects the image quality. Following previous work (Lyu et al., 2023; Kawar et al., 2023), the results are averaged over 150 samples, each with 10 editing prompts. S2Edit achieves the best results in all metrics, showing its superiority in editing quality.
In Fig. 7, we investigate the trade-off between prompt alignment and identity preservation measured by LPIPS, i.e., feature similarity between $\mathcal{I}$ and $\mathcal{I}^{\ast}$, and CLIP score (Hessel et al., 2021), i.e., feature similarity between $\mathcal{P}^{\ast}$ and $\mathcal{I}^{\ast}$. LPIPS and CLIP score are competing metrics, since increasing the degree to which the edited images aligned with target prompts will reduce their similarity with the source images. Still, Fig. 7 shows that for a broad range of LPIPS values, our method has better prompt alignment for the same amount of image changes compared to other methods.

### v2

@@ 1 Introduction
Recent years have witnessed remarkable progress in the field of generative models, with techniques such as Generative Adversarial Networks (GANs) (goodfellow2014generative; karras2019style; karras2020analyzing; karras2021alias) and Diffusion Models (sohl2015deep; saharia2022photorealistic; unCLIP; rombach2022high) capable of synthesizing realistic images. These advancements have fueled significant interest in the realm of image editing, a highly practical and versatile domain. Among its various settings, text-guided editing stands out for its user-friendly interface, as it accepts natural language as the input.
@@
Early methods in this field often train GANs from scratch on paired image-text data (GANEdit1; GANEdit2; GANEdit3), requiring expensive annotations. Later approaches explore the latent space of pre-trained models like CLIP (radford2021learning), which align text and image features. However, GANs are still trained on relatively small datasets, resulting in a latent space with limited capacity. They sometimes struggle to represent real-world objects at the tails of distributions faithfully, and fail to generalize to unseen text prompts. Recent advances in large-scale pre-trained text-to-image diffusion models (saharia2022photorealistic; rombach2022high) offer alternative solutions to this task. Thanks to the large scale pre-trained data, these methods exhibit a significantly stronger capacity in representing diverse objects from different domains, and are generally more robust to text prompt variations. Nevertheless, demonstrations of these editing methods’ performance typically involve marked changes in art styles, object shapes, motions, and precise editing of real-world images such as faces remains under-explored. When guided by detailed prompts, these methods frequently introduce severe artifacts or unrealistic styles. Moreover, they sometimes fail to edit the image utterly, with object identity information changed or the target attribute indicated in the prompt missing in the output. Essentially, current methods struggle to balance between identity information preservation and target prompt alignment (SINE; kawar2023imagic; brooks2023instructpix2pix), and significant challenges still remain in editing real-world images both accurately and faithfully.
@@
In response to these challenges, we propose S2Edit, a two-stage text-guided image editing method. The first stage involves fine-tuning a pre-trained diffusion model to reconstruct the input image conditioned on a text prompt, which helps capture the object identity. To apply target edits while keeping the original identity information intact, we insert a learnable identity token into the prompt to capture nuanced identity features. However, plain fine-tuning lacks the control of the semantic meaning or spatial emphasis of the identity token, which is critical for precise image editing. We observe that the learned identity token often contains information on attributes targeted for modification, which impedes the later editing process. To disentangle the identity token from target attributes, we apply semantic control by enforcing an orthogonality constraint between the identity token and the prompt in the textual feature space during fine-tuning. Moreover, the identity token may learn to represent identity-irrelevant image regions, undermining the preservation of identity information. Therefore, we apply spatial control during fine-tuning by manipulating the cross-attention maps of the identity token with an object mask. In the second stage of S2Edit, we augment the cross-attention injection strategy from Prompt-to-Prompt (hertz2022prompt) with spatial control, and employ it for editing with our identity token. With its high-quality editing results, S2Edit outperforms state-of-the-art methods both quantitatively and qualitatively in evaluations, and garners at least 45% more preference from the participants in a user study.
@@ 2 Related Work
Diffusion-based Image Editing. First introduced by Sohl-Dickstein et al. (sohl2015deep), diffusion models have recently achieved state-of-the-art image synthesis quality (ho2020denoising; dhariwal2021diffusion; nichol2021improved; saharia2022photorealistic; rombach2022high). Based on these models, image editing works perform tasks like inpainting (song2020score; paint-by-example), stroke-based editing (meng2021sdedit; BlendedDMEdit; BlendedLDMEdit), and text-guided editing (hertz2022prompt; PnP-Diffusion; parmar2023zero). In particular, text-guided editing garners considerable interest due to its friendly interface and substantial control capabilities. Studies in this domain enable rich text guidance by manipulating attention maps (hertz2022prompt; PnP-Diffusion; parmar2023zero), text prompts (nichol2021glide; mokady2023null; brooks2023instructpix2pix; kawar2023imagic; pan2023effective; PromptTuningInversion), and feature space (kim2022diffusionclip; bansal2023universal). Prompt-to-Prompt (hertz2022prompt) investigates the relation between prompt tokens and spatial layout of images, and edits images by manipulating the cross-attention layers. InstructPix2Pix (brooks2023instructpix2pix) fine-tunes a pre-trained model on synthesized paired images and editing instructions. Despite impressive results, Prompt-to-Prompt sometimes fails to balance identity preservation and desired editing effects, and its performance often depends on the quality of the given prompt. InstructPix2Pix partially addresses the second issue with more straightforward instruction guidance, yet the first issue remains as it does not provide identity-related control.
@@
Personalized Image Synthesis and Manipulation. To preserve the fidelity of the original image, recent works in the image diffusion domain propose to insert learnable tokens and fine-tune pre-trained text-to-image models such that identity information can be learned and re-combined with new contexts (gal2022image; kumari2023multi; ruiz2023dreambooth; shi2023instantbooth). DreamBooth (ruiz2023dreambooth) learns to encode object identity in a unique token by fine-tuning diffusion models on several images containing the same object. SINE (SINE) investigates the overfitting problem when fine-tuning involves only one image, and proposes to distill the knowledge of the fine-tuned model into the original model. Instead of fine-tuning the entire model, Custom Diffusion (kumari2023multi) only optimizes the Key and Value projection layers in cross-attention, and is able to compose multiple concepts via joint training on images with various objects. Although personalized fine-tuning helps preserve object identity, we observe a lack of semantic and spatial control over the learned identity token in existing methods, leading to failure modes such as entangled concepts when applied to detailed editing tasks.
@@
GAN-based Image Editing. One line of work on GAN attempts to discover meaningful feature space directions (harkonen2020ganspace; shen2020interfacegan; chen2022exploring; wu2021stylespace) that correspond to different semantic changes of images. Some other methods invert images to the latent space of pre-trained GANs (richardson2021encoding; tov2021designing; wang2022high; lyu2023deltaedit; DM-GAN-Edit) and leverage the aforementioned semantic directions for editing. This direction is further spurred by the development of large vision-language models such as CLIP (radford2021learning). DeltaEdit (lyu2023deltaedit) proposes a CLIP delta space where visual and textual features from CLIP are better aligned, and trains a mapping network that connects the delta space to StyleGAN’s (karras2020analyzing) latent space to enable text-guided editing. However, compared to state-of-the-art diffusion models, GANs, which are trained on smaller datasets, have a more limited expressibility in their latent space.
@@ 3 Method
Figure 2: S2Edit overview. Left: Given a source image and a text prompt, we insert a learnable token [I] into the text prompt and fine-tune a pre-trained text-to-image diffusion model to learn the identity information. To obtain a disentangled identity token, we apply an orthogonality constraint in the text embedding space via a semantic loss $L_{semantic}$ and force [I] to only represent the object of interest with masked cross-attention. Right: With [I] learned, we freeze the fine-tuned model and perform Null-text Inversion (mokady2023null) to get an initial noise map, then denoise it conditioned on the target prompt to generate the editing result.
@@ 3.1 Identity Fine-tuning
Instead of optimizing the prompt in the text space, we adapt the special token introduced in DreamBooth (ruiz2023dreambooth) to our method by constructing enhanced prompts $\widetilde{\mathcal{P}}$ and $\widetilde{\mathcal{P}^{\ast}}$ with a unique identity token [I] (e.g. “a [I] man with beard”), which encodes rich identity information and provides clear guidance to the generation process. Different from DreamBooth, which aims to extract appearance of a subject from multiple images for recontextualization tasks, we learn the identity information from one image to ensure that subtle details can be flawlessly preserved. We fine-tune the diffusion model and its text encoder with $\mathcal{I}$ and the enhanced prompt $\widetilde{\mathcal{P}}$ to learn the identity information, which is encoded into both the embedding of [I] and the model weights.
@@ 3.2 Precise Semantic and Spatial Control
where $\mathcal{A}_{\text{[I]}}$ is the original query-key cross-attention map of [I], $\mathcal{A}^{\ast}_{\text{[I]}}$ is its masked version, $M_{obj}$ is the binary mask of the object of interest, and $\odot$ is element-wise product. To obtain $M_{obj}$, we take the cross-attention map of the word that describes the object of interest (e.g., “lady”, “cat”) and binarize it, thus requiring no mask annotation. Compared to other text-guided image editing techniques, such as DeltaEdit, Imagic, SINE, and Prompt Tuning Inversion (lyu2023deltaedit; kawar2023imagic; SINE; PromptTuningInversion), our approach imposes no additional demands on users, apart from specifying the object within the source prompt.
@@ 4.1 Experimental Setup
Implementation Details. While S2Edit can be applied to any text-to-image diffusion model, we implement it based on Stable Diffusion v1-4 (rombach2022high) due to its public availability and importance in the literature. During identity fine-tuning, we use the AdamW (adamw) optimizer with a base learning rate of $2\times 10^{-6}$. The semantic loss weight $\lambda$ is set to $0.1$. We fine-tune the model for 200 steps to avoid overfitting. For the DDIM scheduler, the number of diffusion steps $T$ is set to 50, and the other hyperparameters are the same as Stable Diffusion (rombach2022high). All experiments are conducted on one NVIDIA A100 GPU, where fine-tuning takes around 95 seconds, Null-text inversion process takes 113 seconds, and the inference time is 9 seconds. For the same source prompt and input image, our method requires fine-tuning and inversion to be run once, then only needs 9 seconds to generate edited results with different target prompts.
Datasets. To verify the editing performance and generalization ability of our proposed method, we conduct extensive experiments on images of diverse objects. Given the higher sensitivity of humans to detailed alterations and unnatural features in human faces, face editing serves as an ideal experiment for editing performance evaluation. Hence, we categorize source images into face and non-face domains, showcasing more results in face editing following (lyu2023deltaedit). For the human face domain, we use images from FFHQ (karras2019style) and CelebA (celeba). For non-face domains, we provide results on AFHQ (afhq), LSUN (lsun) cat and church images.
Comparisons. We compare S2Edit with several state-of-the-art diffusion-based text-guided image editing methods: Null-text Inversion (mokady2023null) combined with Prompt-to-Prompt (hertz2022prompt), InstructPix2Pix (brooks2023instructpix2pix), and SINE (SINE). InstructPix2Pix accepts editing instructions as its prompt inputs, such as “Make her smile”. In addition, we compare with a state-of-the-art GAN-based method DeltaEdit (lyu2023deltaedit), which takes prompts inputs containing the word “face”, e.g., “Face with smile”.
@@ 4.2 Qualitative Evaluation
Fig. 3 shows the qualitative comparison between our method and four state-of-the-art methods on face editing. Our method faithfully preserves the original identities while accurately performing the desired edits. DeltaEdit (lyu2023deltaedit) follows the editing prompts well, but fails to preserve the identity of source images due to the limited capacity of GANs’ latent space. For example, the hairstyle is changed in the top row, and the expression drifts in the fourth row. InsturctPix2Pix (brooks2023instructpix2pix) makes the edits but fails to locate the desired editing area, e.g., removing only part of the glasses in the third row, and colorizing the glasses which are undesired in the last two rows. Except for the eyeglass removal failure, Null-text Inversion (mokady2023null) struggles to balance identity preservation and prompt alignment. For instance, the glasses are not removed in the third row, and identity drifts in the bottom row. SINE (SINE) mostly produces unrealistic results with severe artifacts. In Fig. 4, we experiment with different edits for the same human, showing versatile editing capability of S2Edit. Only the specified attribute of interest is manipulated, while irrelevant features are unaltered, which proves that S2Edit learns disentangled identity tokens. Overall, S2Edit achieves the most accurate editing results on human faces which requires precise control of localized details.
@@ 4.3 Quantitative Evaluation
We report common metrics of all methods including FID (FID), LPIPS (lpips), and PSNR in Tab. 1. LPIPS and PSNR measure identity preservation, and FID reflects the image quality. Following previous work (lyu2023deltaedit; kawar2023imagic), the results are averaged over 150 samples, each with 10 editing prompts. S2Edit achieves the best results in all metrics, showing its superiority in editing quality.
In Fig. 7, we investigate the trade-off between prompt alignment and identity preservation measured by LPIPS, i.e., feature similarity between $\mathcal{I}$ and $\mathcal{I}^{\ast}$, and CLIP score (clipscore), i.e., feature similarity between $\mathcal{P}^{\ast}$ and $\mathcal{I}^{\ast}$. LPIPS and CLIP score are competing metrics, since increasing the degree to which the edited images aligned with target prompts will reduce their similarity with the source images. Still, Fig. 7 shows that for a broad range of LPIPS values, our method has better prompt alignment for the same amount of image changes compared to other methods.